The area of the rectangle is 100 square cm. If a square built using a diagonal of this rectangle as one of its sides, the square has area 700 square cm. What is the perimeter of the rectangle?
Answer by mananth(16949)

The area of the rectangle is 100 square cm. If a square built using a diagonal of this rectangle as one of its sides, the square has area 700 square cm. What is the perimeter of the rectangle?
let length be x
width be y
xy=100
..
the diagonal of the rectangle = sqrt(x^2+y^2)
this is the side of the square
Area of this square = sqrt(x^2+y^2)*sqrt(x^2+y^2)
= x^2+y^2
..
x^2+y^2=700
x^2+2xY+y^2= 700+2xy
(x+y)^2= 700+200
(x+y)^2= 900
(x+y)= 30
(x+100/x)=30
x^2+100 /x =30
x^2+100=30x
x^2-30x+100=0
The roots of the equation x1 , x2
x1= 30 +sqrt(900-400) /2
x1=26.18 cm
x2= 30 -sqrt(900-400) /2
x2=3.82 cm
..
perimeter of rectangle = (2*3.82)+(2*25.18)
=60 cm